OpenStudy (anonymous):

Which Value is a solution to (Look in reply bar) A. 1 B. 4 C. 6 D. 8

OpenStudy (anonymous):

\[J - 6 \ge 1\]

OpenStudy (anonymous):

\(j-6\geq 1\) and so \(j\geq 7\) pick the one greater than or equal to seven

OpenStudy (anonymous):

thank you ^^